Hi I am really having a hard time with this problem. We are learning about polynomials and the problem is:
Solve for X
the solution will be AN EXPRESSION WITH a and or b in it
x+2a=3x-a
answer will be 3a/2

Answer by Nate(3500)

x + 2a = 3x - a
x + 2a + a = 3x - a + a
x + 3a = 3x
x - x + 3a = 3x - x
3a = 2x
3a/2 = x